a length of a rectangle is 5 more than the width and the perimeter of a rectangle is 6 times the width.find the width length and perimeter of the rectangle.

Answer by jessica43(140)


To solve this problem you need to set up two equations using what you know. First, you know the length of the rectangle is 5 more than the width: L = 5 + W. Second, you know the perimeter of the rectangle is 6 times the width: L + L + W + W = 6W, or 2L + 2W = 6W.
Now plug in the first equation into the second equation, substituting the L with the (5 + w):
2L + 2W = 6W
2(5 + W) + 2W = 6W
10 + 2W + 2W = 6W
10 + 4W = 6W
10 = 2W
5 = W
So the width is 5. Now plug that into the first equation:
L = 5 + W
L = 5 + 5
L = 10. So using the length and width you can find the perimeter:
2L + 2W = P
2(10) + 2(5) = P
20 + 10 = P
30 = P. The perimeter is 30.